Given the demand function P = 20 – 5Q, find the price elasticity of demand when price of the
commodity is 5 Birr per unit. Mention if the demand is price elastic or inelastic at this point.
1
Expert's answer
2021-07-14T12:52:15-0400

"PED ="


"=\frac{ Q_{2} - Q_{1} }{ (Q_{2} + Q_{1})/2 }*100"


"\frac{ 3_{} - 3.8_{} }{ (3_{} + 3.8_{})/2 }*100=-23.53"


"=\frac{ P_{2} - P_{1} }{ (P_{2} + P_{1})/2 }*100"


"\frac{ 5_{} - {1} }{ (5_{} + 1_{})/2 }*100=133.33"


"PED"


"=\frac{−23.53 }{133.33 }"


"=0.18"

The demand price is inelastic